# Evaluating Compiler Optimization Impacts on zkVM Performance

### v1

@@ Abstract.
We present the first systematic study of the impact of compiler optimizations on zkVMs. We evaluate 64 LLVM passes, six standard optimization levels, and an unoptimized baseline across 58 benchmarks on two RISC-V–based zkVMs (RISC Zero and SP1). While standard LLVM optimization levels do improve zkVM performance (over 40%), their impact is far smaller than on traditional CPUs, since their decisions rely on hardware features rather than proof constraints. Guided by a fine-grained pass-level analysis, we slightly refine a small set of LLVM passes to be zkVM-aware, improving zkVM execution time by up to 45% (average +4.6% on RISC Zero, +1% on SP1) and achieving consistent proving-time gains. Our work highlights the potential of compiler-level optimizations for zkVM performance and opens new direction for zkVM-specific passes, backends, and superoptimizers.
@@ 1. Introduction
Despite this progress, performance remains the central hurdle and a competitive factor among zkVM vendors (Guibas, 2024; Ethereum Foundation, 2025). zkVMs’ architecture adopts a cost model fundamentally different from traditional CPUs, where every instruction becomes a set of constraints to be proved. Yet, compilers still optimize for hardware features that are not relevant to zkVMs, such as caches, out-of-order execution, and instruction-level parallelism. All these raise the following question: How do compiler-level techniques (i.e., optimizations) influence zkVM performance, and to what extent?
This open question motivates this study, which aims to investigate how traditional optimizations influence zkVM performance. To do so, we seek answers for the following three research questions:
@@ 3. RQ3
[…]
@@
  • We present the first systematic study of how traditional compiler optimizations (both individual passes and their combinations) affect zkVM performance, and compare the results to traditional CPUs.
  • We analyze selected optimizations with notable performance effects on zkVMs, extract insights about the underlying root causes, and use them to refine these optimizations for better zkVM performance.
  • We enumerate the implications of our findings, and discuss potential future directions on improving zkVM performance via compiler-level techniques.
Summary of findings. Some of our representative findings are: (1) default optimization levels (e.g., -O3) improve both zkVM execution and proving time by over 40%; (2) however, these gains are significantly smaller than on traditional CPUs, as many individual passes exhibit opposing effects, mainly due to wrong decision heuristics for inlining, loop unrolling, branch elimination, or mismatches in the cost models for arithmetic and memory operations; and (3) autotuning proves to be an effective approach for performance-critical zkVM programs (up to $2.2\times$ speedup compared to -O3).
Implications. Building on our findings, we identify root causes of inefficiencies in certain passes and implement slight modifications to an existing compiler toolchain (i.e., LLVM). The -O3 level in our modified LLVM outperforms the original LLVM’s -O3 in most benchmarks, achieving up to 45% faster zkVM execution (avg. +4% on RISC Zero; +1% on SP1) while also delivering consistent improvements in proving time.
@@ 2. Background
SNARKs target circuit satisfiability (Parno et al., 2013), representing statements as arithmetic circuits. Early ZK development used low-level ZK DSLs (Bellés-Muñoz et al., 2023), where a compiler arithmetizes the circuit into a concrete constraint system (e.g., PLONK-style (Gabizon et al., 2019)) and produces a witness generator. Then, the prover uses the constraints and witness to produce a proof. Writing low-level ZK circuits is not only error-prone (Chaliasos et al., 2024a), but it also requires significant expertise. Zero-knowledge virtual machines (zkVMs) (Ben-Sasson et al., 2014) raise the abstraction by compiling source code (e.g., Rust) to a target ISA (often RISC-V (Waterman et al., 2016)(Labs, 2024b; Bruestle and Gafni, 2023). Then, an emulator executes the binary and records an execution trace in the proper arithmetization, and the back-end produces a proof. In this context, developers use well-established compilation toolchains (e.g., LLVM), and the zkVM exposes precompiles for heavy primitives (e.g., SHA-2) to replace thousands of generic instructions with optimized circuits. To make zkVMs practical for computation-heavy applications (e.g., ZK-Rollups (Barry Whitehat, 2018)), much of the literature focuses on improving both execution and proof generation (Arun et al., 2023; Setty et al., 2023; Goldberg et al., 2021; Labs, 2024b; Bruestle and Gafni, 2023). Figure 1 contrasts native execution, ZK DSLs, and zkVMs. For an in-depth introduction of ZKPs, we refer the reader to (Thaler et al., 2022).
@@
zkVMs vs. traditional architectures. When generating the execution trace and proving it on a zkVM, the cost model differs significantly from traditional CPU architectures, such as x86 or ARM. Most instructions have near-uniform cost, and there is no instruction latency due to cache hits/misses and other micro-architectural effects. Memory access is relatively cheaper on zkVMs, but page-ins and page-outs can be significantly more expensive. Further, there is no parallelism or multithreading, and floating-point operations are not supported natively, so they need to be emulated, which makes them extremely costly. We detail a full list of differences in Appendix A. Notably, zkVM execution and proving are orders of magnitude slower than native execution (milliseconds vs. seconds to hours, cf. Figure 14 Appendix A).
@@ 3.1. Metrics
We evaluate the impact of optimizations on zkVMs using three zkVM metrics: cycle count, zkVM execution time, and proving time. Cycle count is the total cycles of the program run on the zkVM (also known as the guest program), and is analogous to CPU cycles. zkVM execution time is the wall-clock time for the executor to replay the guest and produce the full execution trace. Proving time is the wall-clock time for the prover to generate a proof and is typically larger than execution time. For native runs (RQ3), we measure native execution time to contrast optimization effects on zkVMs and traditional architectures.
@@ 3.2. Benchmark Selection
We combine standard benchmarks along with zkVM-focused workloads to cover both large, optimization-sensitive programs and common zkVM payloads. In total, our benchmark suite consists of 58 programs written in Rust and C/C++. We use PolyBench (Pouchet, 2010; Ferrer, 2020), NPB (Division, 2024; Martins et al., 2025), and SPEC CPU 2017 (Bucek et al., 2018), which include compute and memory-intensive programs. From industry benchmark suites, we include the a16z crypto zkVM, the Succinct Labs, and the RSP benchmarks (a16z crypto, 2024; Labs, 2024d, a). These capture typical cryptography-heavy workloads. We further add targeted programs, such regex-match, sha256, and loop-sum, to exercise loop-heavy, memory-heavy, and math/crypto patterns frequently seen in zkVMs. For the complete list of the benchmarks see Appendix B.
@@ 3.3. Analyzing Optimizations
To answer the RQs outlined in Section 1, we first measure each pass in isolation on RISC Zero and SP1 across our benchmark suite, collecting cycle count, executor time, and proving time (Section 4.1) and quantify the performance losses/gains compared to the baseline. Next, we explore optimization combinations via genetic autotuning with OpenTuner (Ansel et al., 2014) (Section 4.2). Then, for RQ3, we run the same optimization profiles on traditional CPU (i.e., x86) and compare the speedups with those obtained from zkVMs to pinpoint divergences caused by the different cost models (Section 4.3).
Setup. x86 and zkVM execution experiments run on an AMD EPYC 7742 with 500GB of RAM on Ubuntu 22.04.5. Proving runs on an AMD EPYC 7542 with 32GB RAM on Ubuntu 24.04.2, equipped with an NVIDIA GeForce RTX 4090 (24GB VRAM). We use SP1 4.1.4 (moongate-server 4.1.0), RISC Zero 1.2.4, Clang 19.1.7, LLVM 19.1.7, and Rust 1.85.0. For both zkVMs we used their default STARK-based proof systems.
@@ 3.4. Threats to Validity
Internal validity. Proving time on SP1 exhibits higher variance than on RISC Zero, partly due to the closed-source prover accessed via RPC and the associated network overhead. To mitigate this, We use at least 10 samples per configuration and increase the sample size for noisy cases. We aim to study broad optimization patterns rather than absolute performance. Cycle count is strongly but not perfectly correlated with proving time, so in edge cases it can over- or underestimate actual costs.
@@
[…]
@@ 4.1. RQ1: Impact of Individual LLVM Passes
Figure 3 presents the top 25 LLVM passes with the highest average impact (either positive or negative) across all benchmarks, zkVMs, and metrics. Each result reflects the performance difference relative to an unoptimized baseline with no passes applied. The remaining 39 passes have minimal impact and are omitted for brevity.
[…]
Inlining is beneficial for both zkVMs: the inline and always-inline passes lead to performance gains of 28.4% and 5.7% on RISC Zero, and 19.3% and 5.3% on SP1, respectively. Other passes, such as instcombine or sroa, perform well on some benchmarks, while degrade performance on others. Both zkVMs exhibit similar trends, with no pass consistently having opposite effects on RISC Zero and SP1.
@@
We compare cycle count with execution and proving times and find a strong linear correlation between them. On both zkVMs, the Pearson and Spearman correlation coefficients exceed 0.98. This confirms that increases in cycle count translate directly into longer execution and proving times. zkVM execution time and proving time show a similar strong linear correlation, with Pearson and Spearman scores above 0.98.
Figure 4. Counting the cases where each optimization pass leads to (1) severe performance gain ($\geq 5$%), moderate performance gain (between $2\%$ and $5\%$), moderate performance loss (between -2% and -5%), and severe performance loss ($\leq-5$%).
@@
[…]
@@
Categories of effects. For completeness, Figure 4 shows the number of benchmarks where each optimization pass causes: (1) moderate performance loss (-5% to -2%), (2) severe loss ($\leq-5\%$), (3) moderate gain (2% to 5%), or (4) severe gain ($\geq 5\%$). Inlining consistently improves both execution and proving times, though a few benchmarks still see execution slowdowns (more details in Section 5). In contrast, loop optimizations (e.g., loop-extract, licm, loop-deletion, loop-unroll) often harm zkVM performance, especially on RISC Zero. This is because loop optimizations in LLVM operate in Loop Closed SSA (LCSSA) form, where values defined in a loop but used outside it are routed through phi nodes at the loop exit. This transformation often requires recomputing addresses (e.g., via getelementptr) or adding load/store instructions, especially when iterating over arrays, leading to extra memory operations that harm zkVM performance (Section 5). Finally, some passes, such as instcombine on RISC Zero execution, show a balanced impact, with similar numbers of benchmarks improving and regressing. This suggests that certain rewrites in instcombine may hurt zkVM performance and should be adjusted to better align with zkVMs’ model (Figure 2(a)).
@@ 4.2. RQ2: Combinations of Optimizations
Standard optimization levels. Figure 5 shows the impact of standard optimization levels (-Ox flags) on execution and proving time across both zkVMs. For brevity, the effects on cycle count are omitted. Excluding -O0, no -Ox level negatively impacts performance relative to the baseline. In contrast, -O0 (i.e., Rust’s MIR optimizations) causes regressions in 19 programs on RISC Zero and 9 on SP1. On average, standard optimization levels improve execution time by 60.5% on RISC Zero and 47.3% on SP1, and proving time by 55.5% and 51.1%, respectively. The -O3 level consistently achieves the highest average performance gains, while -Oz (reducing binary size) shows the lowest.
@@
Autotuning. Overall, the existing -Ox flags have a positive effect on both zkVM execution and proving time. However, can we identify sequence of passes that outperform them? To explore this, we employ OpenTuner (Ansel et al., 2014), a genetic-based tuning framework that finds an optimal combination of compilation flags guided by a specific fitness function. Although zkVM execution and proving time are the ultimate metrics of interest, they are too expensive to evaluate repeatedly during autotuning. Instead, we use cycle count as a proxy, based on its strong linear correlation with both execution and proving established in RQ1 (Section 4.1). This choice offers two key advantages: (1) cycle count avoids measurement noise, and (2) it is fast to compute ($\sim$1 second on average).
[…]
@@
[…]

### v2

@@ Abstract.
We present the first systematic study of the impact of compiler optimizations on zkVMs. We evaluate 64 LLVM passes, six standard optimization levels, and an unoptimized baseline across 58 benchmarks on two RISC-V–based zkVMs (RISC Zero and SP1). While standard LLVM optimization levels do improve zkVM performance (over 40%), their impact is far smaller than on traditional CPUs, since their decisions rely on hardware features rather than proof constraints. Guided by a performance impact analysis, we slightly refine a small set of LLVM passes to be zkVM-aware, improving zkVM execution time by up to 45% (average +4.6% on RISC Zero, +1% on SP1) and achieving consistent proving-time gains. Our work highlights the potential of compiler-level optimizations for zkVM performance and opens new directions for zkVM-specific passes, backends, and superoptimizers.
@@ 1. Introduction
Despite this progress, performance remains the central hurdle and a competitive factor among zkVM vendors (Guibas, 2024; Ethereum Foundation, 2025). zkVMs’ architecture adopts an execution model fundamentally different from traditional CPUs, where every instruction becomes a set of constraints to be proved. Yet, compilers still optimize for hardware features that are not relevant to zkVMs, such as caches, out-of-order execution, and instruction-level parallelism. All these raise the following question: How do compiler-level techniques (i.e., optimizations) influence zkVM performance, and to what extent?
This open question motivates this study, which aims to investigate how existing optimizations influence zkVM performance and identify concrete opportunities for improvement. We seek answers for the following three research questions:
@@ 3. RQ3
To answer these research questions, we benchmark 58 programs on two zkVMs, RISC Zero (Bruestle and Gafni, 2023) and SP1 (Labs, 2024b), across 71 optimization profiles that include 64 individual LLVM passes, six default optimization levels (e.g., -O3), and an unoptimized baseline. The selected zkVMs are the most mature and widely used ones, targeting a conventional ISA that leverages well-established toolchains (Zero, 2025b; Labs, 2025), such as LLVM. Our benchmarks cover a wide range of programs, including cryptographic primitives and mathematical computations. The effects of optimizations are quantified through three metrics: cycle count, zkVM execution time, and proving time.
@@
  • We present the first systematic study of how traditional compiler optimizations (both individual passes and their combinations) affect zkVM performance, and compare the results to traditional CPUs (Section 4).
  • We identify and empirically validate key cost components that drive zkVM performance, that is, dynamic instruction count, and paging cycles. (Section 5.1).
  • We analyze selected optimizations based on our cost components, and derive four optimization principles, which we validate empirically. We then apply these principles to refine existing passes for better zkVM performance. (Sections 5.2 and 6.1).
  • We enumerate the implications of our findings, and discuss potential future directions on improving zkVM performance (Section 6.2).
Summary of findings. Some of our representative findings are: (1) default optimization levels (e.g., -O3) improve both zkVM execution and proving time by over 40%; (2) however, these gains are significantly smaller than on traditional CPUs, as many individual passes exhibit opposing effects, mainly due to wrong decision heuristics for inlining, or loop unrolling. These heuristics rely on exploiting hardware-specific features (e.g., cache locality) rather than minimizing the number of the executed instructions; and (3) autotuning proves to be an effective approach for performance-critical zkVM programs (up to $2.2\times$ speedup compared to -O3).
Implications. We identify root causes of inefficiencies in certain passes and implement slight modifications to an existing compiler toolchain (i.e., LLVM). The -O3 level in our modified LLVM outperforms the original LLVM’s -O3 in most benchmarks, achieving up to 45% faster zkVM execution (avg. +4% on RISC Zero; +1% on SP1) while also delivering consistent improvements in proving time. Since zkVM execution and proving are orders of magnitude more expensive than execution on traditional CPUs, even small percentage improvements can translate into minutes of savings (Section 2).
@@ 2. Background
SNARKs target circuit satisfiability (Parno et al., 2013), representing statements as arithmetic circuits. Early ZK development used low-level ZK DSLs (Bellés-Muñoz et al., 2023), where a compiler arithmetizes the circuit into a concrete constraint system (e.g., PLONK-style (Gabizon et al., 2019)) and produces a witness generator. Then, the prover uses the constraints and witness to produce a proof. Writing low-level ZK circuits is not only error-prone (Chaliasos et al., 2024a), but it also requires significant expertise. Zero-knowledge virtual machines (zkVMs) (Ben-Sasson et al., 2014) raise the abstraction by compiling source code to a target ISA (often RISC-V (Waterman et al., 2016)(Labs, 2024b; Bruestle and Gafni, 2023). Then, the zkVM’s emulator executes the binary and records an execution trace in the proper arithmetization, and the back-end produces a proof. Developers use well-established compilation toolchains (e.g., LLVM), and the zkVM exposes precompiles for heavy primitives (e.g., SHA-2) to replace thousands of instructions with optimized circuits. To make zkVMs practical for computation-heavy applications (e.g., ZK-Rollups (Barry Whitehat, 2018)), much of the literature focuses on improving both execution and proof generation (Arun et al., 2023; Setty et al., 2023; Goldberg et al., 2021; Labs, 2024b; Bruestle and Gafni, 2023). Figure 1 contrasts native execution, ZK DSLs, and zkVMs. For an in-depth introduction of ZKPs, we refer the reader to (Thaler et al., 2022).
@@
zkVMs vs. traditional architectures. When generating the execution trace and proving it on a zkVM, the execution model differs significantly from traditional CPU architectures, such as x86 or ARM. Most instructions have near-uniform cost, and there is no instruction latency due to cache hits/misses and other micro-architectural effects. Memory access is relatively cheaper on zkVMs (cache misses are not penalized), but page-ins and page-outs can be significantly more expensive. Further, there is no parallelism or multithreading, and floating-point operations are not supported natively, so they need to be emulated, which makes them extremely costly. We detail a full list of differences in Appendix A. Notably, zkVM execution and proving are orders of magnitude slower than native execution (milliseconds vs. seconds to hours, cf. Figure 15 Appendix A).
@@ 3.1. Metrics
We evaluate the impact of optimizations on zkVMs using three zkVM metrics: cycle count, zkVM execution time, and proving time. Cycle count is the total cycles of the program run on the zkVM (also known as the guest program), and is analogous to CPU cycles. It captures only the cost of program execution and proving, and does not include other one-time setup or initialization overheads, such as prover initialization. zkVM execution time is the wall-clock time for the executor to replay the guest and produce the full execution trace. Proving time is the wall-clock time for the prover to generate a proof and is typically larger than execution time. For native runs (RQ3), we measure native execution time to contrast optimization effects on zkVMs and traditional architectures.
@@ 3.2. Benchmark Selection
We combine standard benchmarks along with zkVM-focused workloads to cover both large, optimization-sensitive programs and common zkVM payloads. In total, our benchmark suite consists of 58 programs written in Rust and C/C++. We use PolyBench (Pouchet, 2010; Ferrer, 2020), NPB (Division, 2024; Martins et al., 2025), and SPEC CPU 2017 (Bucek et al., 2018), which include compute and memory-intensive programs. From industry benchmark suites, we include the a16z crypto zkVM, the Succinct Labs, and the RSP benchmarks (a16z crypto, 2024; Labs, 2024d, a). These capture typical cryptography-heavy workloads. We further add targeted programs, such as regex-match, sha256, and loop-sum, to exercise loop-heavy, memory-heavy, and math/crypto patterns frequently seen in zkVMs. For the complete list of the benchmarks see Appendix B.
@@ 3.3. Analyzing Optimizations
To answer the RQs outlined in Section 1, we first measure each pass in isolation on RISC Zero and SP1 across our benchmark suite, collecting cycle count, executor time, and proving time (Section 4.1) and quantify the performance losses/gains compared to the baseline. Next, we explore optimization combinations via genetic autotuning with OpenTuner (Ansel et al., 2014) (Section 4.2). Then, for RQ3, we run the same optimization profiles on traditional CPU (i.e., x86) and compare the speedups with those obtained from zkVMs to pinpoint opportunities for improvement and understand which optimizations are worth revisiting or redesigning in the zkVM context (Section 4.3).
Setup. x86 and zkVM execution experiments run on an AMD EPYC 7742 with 500GB of RAM on Ubuntu 22.04.5. Proving runs on an AMD EPYC 7542 with 32GB RAM on Ubuntu 24.04.2, equipped with an NVIDIA GeForce RTX 4090 (24GB VRAM). The SP1 prover is executed locally via its RPC API. We use SP1 4.1.4 (moongate-server 4.1.0), RISC Zero 1.2.4, Clang 19.1.7, LLVM 19.1.7, and Rust 1.85.0. For both zkVMs we used their default STARK-based proof systems.
@@ 3.4. Threats to Validity
Internal validity. Proving time on SP1 exhibits higher variance than on RISC Zero, partly due to the closed-source prover accessed via RPC and the associated network overhead. To mitigate this, We use at least 10 samples per configuration and increase the sample size for noisy cases.
@@
Construct validity. Several benchmarks use reduced input sizes to keep proving feasible. We evaluate 64 individual LLVM passes essentially in isolation with default parameters. Many additional passes and tune strategies exist, and some benefits only appear in specific combinations. Although we explored combinations via autotuning, the whole phase-ordering space remains intractable.
@@ 4.1. RQ1: Impact of Individual LLVM Passes
Figure 3 presents the top 25 LLVM passes with the highest average impact (either positive or negative) across all benchmarks, zkVMs, and metrics. Each result reflects the performance difference relative to an unoptimized baseline with no passes applied. The remaining 39 passes have minimal impact and are omitted for brevity, as they change execution time and proving time by only 0.9%–1%, on average.
zkVM execution time. Only a few passes, on average, improve zkVM execution time when applied individually. Out of 64 passes, 12 reduce execution time on RISC Zero by at least 1% on average, and 13 do so on SP1. The most harmful pass on both zkVMs is licm, which increases execution time by 11.8% on RISC Zero, and 7.1% on SP1. Section 5 analyzes the root causes of this regression in detail.
Inlining is beneficial for both zkVMs: the inline and always-inline passes lead to performance gains of 28.4% and 5.7% on RISC Zero, and 19.3% and 5.3% on SP1, respectively. Other passes, such as instcombine or sroa, perform well on some benchmarks, while degrading performance on others. Both zkVMs exhibit similar trends, with no pass consistently having opposite effects on RISC Zero and SP1.
@@
Figure 4. Counting the cases where each pass leads to (1) severe performance gain ($\geq 5$%), moderate performance gain (between $2\%$ and $5\%$), moderate performance loss (between -2% and -5%), and severe performance loss ($\leq-5$%).
@@
Execution Proving
zkVM Gain (>2%) Loss (<-2%) Gain (>2%) Loss (<-2%)
@@
Categories of effects. For completeness, Figure 4 shows the number of benchmarks where each pass causes: (1) moderate performance loss (-5% to -2%), (2) severe loss ($\leq-5\%$), (3) moderate gain (2% to 5%), or (4) severe gain ($\geq 5\%$). Inlining consistently improves both execution and proving times, though a few benchmarks still see execution slowdowns (more details in Section 5). In contrast, loop optimizations (e.g., loop-extract, licm, loop-deletion, loop-unroll) often harm zkVM performance, especially on RISC Zero. This is because loop optimizations in LLVM operate in Loop Closed SSA (LCSSA) form, where values defined in a loop but used outside it are routed through phi nodes at the loop exit. This transformation often requires recomputing addresses (e.g., via getelementptr) or adding load/store instructions, especially when iterating over arrays, leading to extra memory operations that harm zkVM performance (Section 5). Finally, some passes, such as instcombine on RISC Zero execution, show a balanced impact, with similar numbers of benchmarks improving and regressing. This suggests that certain rewrites in instcombine may hurt zkVM performance and should be adjusted to better align with zkVMs’ model (Figure 2(a)).
@@ 4.2. RQ2: Combinations of Optimizations
Standard optimization levels. Figure 5 shows the impact of standard optimization levels (-Ox flags) on execution and proving time across both zkVMs, relative to our baseline where no optimizations are applied, including Rust’s MIR optimizations (Section 3.3). For brevity, the effects on cycle count are omitted. Excluding -O0, no -Ox level negatively impacts performance relative to the baseline. In contrast, -O0 (i.e., Rust’s MIR optimizations) causes regressions in 19 programs on RISC Zero and 9 on SP1. On average, standard optimization levels improve execution time by 60.5% on RISC Zero and 47.3% on SP1, and proving time by 55.5% and 51.1%, respectively. The -O3 level consistently achieves the highest average performance gains, while -Oz (reducing binary size) shows the lowest.
@@
Autotuning. Overall, the existing -Ox flags have a positive effect on both zkVM execution and proving time. However, can we identify sequence of passes that outperform them? To explore this, we employ OpenTuner (Ansel et al., 2014), a genetic-based tuning framework that finds an optimal combination of compilation flags guided by a specific fitness function. Although zkVM execution and proving time are the ultimate metrics of interest, they are too expensive to evaluate repeatedly during autotuning. Instead, we use cycle count as a proxy, as it is a natural performance predictor. This choice offers two key advantages: (1) cycle count avoids measurement noise, such as the inherent imprecision of wall-clock timing and (2) it is fast to compute ($\sim$1 second on average).
We autotune each benchmark independently for each zkVM. We consider all LLVM passes as tunable parameters, creating sequences of passes up to a depth of 20. For compiler options that receive integer or enum values (e.g., -inline-threshold), we define appropriate tuning range and use OpenTuner’s parameter types to select the optimal value. For most flags, these ranges are large enough to be effectively unbounded. Our artifact   [^1] https://github.com/thomasgassmann/zkvm-compiler-optimizations includes and documents the tuning ranges for all examined compiler flags. Each OpenTuner run performs 160 iterations. Even with this limited number of iterations, the result of OpenTuner in several benchmarks outperforms -O3. On RISC Zero, 18 out of 58 programs outperform -O3 in terms of cycle count reduction (on average by 6.3%), while on RISC Zero, 20 programs outperform -O3 (on average 14.6% cycle count reduction).
@@
Autotuning specific benchmark suites. In the previous experiment, each benchmark was tuned individually using cycle count as the fitness function. Yet, how does this translate to zkVM execution and proving time? To answer this, we autotune (1) the NPB benchmark (Martins et al., 2025) suite and (2) cryptography benchmarks on both zkVMs, running OpenTuner for a longer time frame (1,600 iterations instead of 160). We select the NPB suite for its large programs, which are more likely to benefit from optimization, and cryptography benchmarks as they best represent typical zkVM workloads.